OpenStudy (anonymous):

Let f(x) = 1/{\sqrt{6x^{2}+6}}. Find f'(x).

OpenStudy (john_es):

Sorry, it should be this way, \[f(x)=(6x^2+6)^{-1/2} \] \[f'(x)=\frac{-1}{2}(6x^2+6)^{-3/2}(12x)=-6x(6x^2+6)^{-3/2}\]

OpenStudy (anonymous):

so the 1 on top doesn't matter?

OpenStudy (john_es):

No, because, \[\frac{1}{x}=x^{-1}\] You can always use this, and differentiate like a power.

OpenStudy (anonymous):

okay thanks again haha :)

OpenStudy (john_es):

You're welcome ;).